From GlueXWiki
Jump to: navigation, search

August 28, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status (Mike, Beni)
  3. FDC status (Vlad, Lubomir)
  4. Electronics (Chris, Nick)
  5. Engineering (Bill)
  6. Commissioning Wiki Page (all)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Sean (NU), Beni, Luke, Chris, Nick, Simon, Fernando, and Lubomir (JLab).


- Beni: testing the latest fADC125 firmware in ROCFDC2. This version has some fixes that Cody did, based on the problems Beni found last week. These tests need DAQ with software trigger. ROCFDC3 has problems: works at the beginning but later has no valid data.

- Lubomir: according to Sasha he made a big progress with the GTP using BCAL north side. There are some minor issues but expect soon to have the BCAL up/down coincidence trigger. When it's ready we have to move the CDC/FDC trigger electronics from the bottom to the top trigger crate. The fan on the bottom trigger crate was replaced and there are no more spare fans; not clear what the problem is with these fan units.

CDC update

- Fernado: working with Nick on the CDC noise from the CAEN supply. Contacted CAEN people: they are trying to be helpful. Grounding of the aluminum ring didn't help. They found two HVBs next to each other, one (C5) is very clean while the other (C6) has huge noise varying from channel to channel. Will use them to troubleshoot the noise. The swapping of HV connectors, or moving HV wires didn't change anything. Tomorrow will continue with "swapping everything" to find the problem.

FDC update

- Lubomir: FDC not operational right now (see chiller problems below). Vlad is comparing the results of the runs before/during/after the magnet test, will report on the next meeting. One TDC channel stopped working. With a scope we see all the channels. Fernando: swap the two signal cables and if the problem is in the same channel contact Ed.


- The heat exchanger was installed on Monday but could not regulate the temperature. According the Bill the problem now is that the input water is too warm. There's a bypass valve and when it's completely open and the LV is OFF it regulates the temperature, but that's not the case when th LV is on. Bill ordered parts for a modification (picture shown) that will fix the problem and probably will be installed tomorrow.

- Even if the chiller doesn't regulate the temperature it gives OK status and the interlock is not engaged, i.e. one can turn the LV and it will overheat. To avoid this, as a first fix, we want to implement in addition an EPICS interlock (discussed already with Hovanes). Later we will need a temperature sensor with a PLS system to interlock the LV.

- Luke finished the cathode scan, latest results shown on the first plot: mean pitch over the area of the cathode. One certainly sees the dips in the seam positions and a little bigger pitch in the middle foil, still there are some irregularities. The latest three measurements were done with the cathode rotated 180deg. One of these was done also with the encoder: the difference between the motor and encoder values has ~1 micron sigma (second plot). Finally, one of the measurement was a repetition of the very first measurement: they are consistent (third plot) within 2-4 microns (the error of finding the edge of the strip with the microscope) but in addition there's some systematic error, a slope as function of the strip number, that we can't explain for now. Luke did a great job with the cathode scan, he's leaving tomorrow.